OpenStudy (anonymous):

what is the value of 2+ (-2/3)^2 ÷ 1/3 ?

OpenStudy (anonymous):

(2+4/9)*3=((18+4)/9)*3=22/3

OpenStudy (anonymous):

u lookig for fraction or approxmiate value?

OpenStudy (anonymous):

that is like multiplying by 3 right?

OpenStudy (anonymous):

yes

OpenStudy (rsadhvika):

2 + (4/9) ÷ 1/3 2 + (4/9) / 3 2 + 4/27 58/27

OpenStudy (anonymous):

my choices are 0 16 3 1/3 and -2

OpenStudy (anonymous):

\[\large 2+ \color {red}{(-2/3)^2} ÷ 1/3 \] \[\large =2+ \frac{4}{9} ÷ \color {red}{1/3} \] \[\large =2+ \frac{4}{9} \times \color {red}{\frac{3}{1}} \] \[\large =2+ \frac{4}{\cancel9^3} \times \frac{\cancel{3}^1}{1} \] \[\large =2+ \frac{4}{3}\] \[\large =2+ 1\frac{1}{3}\] \[\large =3\frac{1}{3}\]